OpenStudy (anonymous):

Check my answer? Question: Write an equation in slope-intercept form for the line that passes through (3, 7) and (7, 4). My answer: Okay so first we use the slope formula to find the slope of the line. y2 - y1/ x2 - x1 = 4 - 7/7 - 3 = - 3/4. So - 3/4 is the slope so now we can use the point slope form. y - y1 = m (x - x1) = y - 7 = - 3/4( x - 3) = y - 7 = - 3/4x + 2.25, now we add seven to both sides. y - 7 + 7 = - 3/4x + 2.25 + 7 = y = - 3/4x + 9.25.

OpenStudy (anonymous):

Check your answer the answer cant be in a decimal

OpenStudy (anonymous):

I have to do it myself to check it. Find slope. \[m =\frac{ y2 - y2 }{ x2 - x1 } = \frac{ 4 - 7 }{ 7 - 3 } = \frac{ -3 }{ 4 }\] So you got the slope right, -3/4. (3, 7) and slope -3/4 Point-Slope form is y - y1 = m (x - x1). y - 7 = -3/4(x - 3) Change to Slope-Intercept form. Distribute. Put into decimal form if its easier. y - 7 = -3/4x + 9/4 or y - 7 = -3/4x + 2.25 Add 7. y = -3/4x + 9.25 If I didn't make any mistakes, I think you are right!

OpenStudy (anonymous):

Thanks for your time :- )

OpenStudy (anonymous):

No problem! (: